# MBSS-T1: Model-Based Self-Supervised Motion Correction for Robust Cardiac T1 Mapping

### v1

@@ Abstract
T1 mapping is a valuable quantitative MRI technique for diagnosing diffuse myocardial diseases. Traditional methods, relying on breath-hold sequences and echo triggering, face challenges with patient compliance and arrhythmias, limiting their effectiveness. Image registration can enable motion-robust T1 mapping, but inherent intensity differences between time points pose a challenge. We introduce MBSS-T1, a self-supervised model for motion correction in cardiac T1 mapping, constrained by physical and anatomical principles. The physical constraints ensure expected signal decay behavior, while the anatomical constraints maintain realistic deformations. The unique combination of these constraints ensures accurate T1 mapping along the longitudinal relaxation axis. MBSS-T1 outperformed baseline deep-learning-based image registration approaches in a 5-fold experiment on a public dataset of 210 patients (STONE sequence) and an internal dataset of 19 patients (MOLLI sequence). MBSS-T1 excelled in model fitting quality ($R^{2}$: 0.974 vs. 0.941, 0.946), anatomical alignment (Dice score: 0.921 vs. 0.984, 0.988), and expert visual quality assessment for the presence of visible motion artifacts (4.33 vs. 3.34, 3.62). MBSS-T1 has the potential to enable motion-robust T1 mapping for a broader range of patients, overcoming challenges such as arrhythmias, and suboptimal compliance, and allowing for free-breathing T1 mapping without requiring large training datasets. Our code and trained models are available at https://github.com/eyalhana/MBSS-T1
@@ 1 Introduction
Quantitative T1 mapping is an advanced MRI technique designed for the precise measurement of intrinsic longitudinal relaxation time in myocardial tissue [26]. This method is increasingly recognized as essential for assessing diffuse myocardial diseases such as inflammation, fibrosis, hypertrophy, and infiltration [26, 28]. “Native” T1 mapping, performed without a paramagnetic contrast agent, effectively detects myocardial conditions, including edema, iron overload, infarcts, and scarring [24]. Furthermore, by allowing for the objective quantification of tissue characteristics, T1 mapping enables longitudinal tracking of changes in the myocardium. This feature is precious for clinical trials, where monitoring the efficacy of interventions over time is critical [24].
@@
Fig. 1 presents the process of generating accurate T1 maps. It begins with capturing a series of images at different time points. Each image captures the myocardium at a distinct point during the longitudinal relaxation process. These images are then analyzed to calculate T1 relaxation times through signal decay model fitting, resulting in detailed T1 maps that reflect the myocardial tissue properties. However, heart motion, respiration, and spontaneous patient movements can cause significant distortions in T1 maps, affecting their reliability and clinical usefulness, and potentially leading to incorrect diagnoses [27]. Echo-triggering is a well-established method for accounting for cardiac motion. Breath-hold sequences, such as the Modified Look-Locker Inversion Recovery (MOLLI) sequence and its variants [23], are often used to minimize respiratory motion artifacts. However, requiring patients to hold their breath presents practical challenges, as not all patients can fully adhere to these instructions. This breath-hold acquisition further limits the number of images that can be captured and makes T1 mapping unsuitable for patients who cannot tolerate breath-holding [23]. Additionally, inadequate echo-triggering due to cardiac arrhythmias can lead to unreliable T1 maps, further compromising diagnostic accuracy.
@@
Despite advancements in classical and deep learning-based methods for motion correction and T1 mapping, significant limitations persist. Classical model-based approaches, while effective, are often slow due to their iterative nature [27, 30]. Deep-learning-based methods, on the other hand, rely heavily on extensively annotated datasets, which are not readily available for all scanners and protocols [8, 31]. Furthermore, these methods typically use either pairwise loss functions based on metrics such as mutual information to guide the registration process [1] or groupwise approaches such as the principal component analysis (PCA), which ignore the signal decay relaxation model [33]. This omission may result in physically unlikely deformations due to changes in image contrast. These limitations hinder the practical applicability and robustness of current methods.
In this work, we present MBSS-T1, a self-supervised, physically, and anatomically constrained deep learning model for simultaneous motion correction and T1 mapping from either breath-hold or free-breathing acquisitions. Our method integrates physical and anatomical insights to correct motion in T1 mapping. The main innovation lies in incorporating the signal decay model within the network’s architecture in a self-supervised framework while utilizing a pretrained segmentation network to produce segmentations that guide the registration and ensure the anatomical correctness of the deformed images. This approach enables the generation of corrected T1 maps across various imaging protocols, eliminating the need for protocol-specific data collection, manual annotations, and extensive training.
This work extends our previous supervised PCMC-T1 model [13], which required a large annotated dataset and extensive training, thereby limiting its generalization capacity across different acquisition settings such as sequence type and the number of frames acquired.
We demonstrated the added value of MBSS-T1 in free-breathing cardiac T1 mapping using a 5-fold experimental setup on a publicly available free-breathing quantitative T1 MRI dataset of 210 patients [4] acquired using the STONE sequence [29] and an in-house dataset of 19 patients acquired using the MOLLI sequence [23] for both free-breathing and breath-hold scans in comparison to baseline methods for deep-learning-based image registration [3, 14] as well as the T1 maps produced on the scanner. We used the fit quality (r-squared), the Dice coefficient, and the Hausdorff distance as quantitative metrics to assess registration accuracy and clinical usability. We further demonstrated the clinical impact through qualitative expert radiologist assessments.
@@ 2 Related Work
Guyader et al. introduced a groupwise image registration technique for quantitative MRI that utilizes a dissimilarity metric derived from an approximated form of total correlation. This method addresses the challenge of aligning multiple images simultaneously without relying on a reference image, thereby reducing registration bias and improving the consistency of quantitative measurements [10]. Tao et al. introduced a robust motion correction method for myocardial T1 and extracellular volume mapping using a PCA-based groupwise image registration technique, which registers all MOLLI frames simultaneously with a total-correlation-based metric [25]. Similarly, Huizinga et al. developed a PCA-based groupwise image registration technique for quantitative MRI (qMRI), including T1 mapping, to correct misalignment without a reference image, thus avoiding registration bias. Their method registers all images to a mean space by minimizing a PCA-derived cost function, assuming that intensity changes in qMRI can be represented by a low-dimensional signal model without needing details of the specific acquisition model [15]. Segmentation-based registration techniques address intensity differences between images by using anatomical structures to guide the registration. El-Rewaidy et al. utilized a non-rigid active shape model-based framework, registering T1-weighted images and computing residual motion by matching manually annotated myocardium contours. This method improves T1 map quality by addressing motion-induced errors in myocardial segments [4].
A key limitation of the aforementioned methods is their failure to account for the signal decay model during registration, resulting in potentially non-physical corrected T1 maps. To address this, model-based approaches have been developed, significantly enhancing T1 mapping accuracy. Xue et al. proposed a fast variational non-rigid registration framework using synthetic image estimation to correct motion in myocardial T1 mapping. Motion-free synthetic images, constructed from initial T1 estimates, guide the registration to avoid motion artifacts [30]. Van Heeswijk et al. introduced a model-based alignment of Look-Locker MRI sequences for myocardial scar quantification, using exponential curve fitting errors as the registration metric [6]. Tilborghs et al. applied a robust motion correction method for cardiac T1 and ECV mapping, integrating data-driven initialization with model-based registration within a non-rigid framework. This method iteratively estimates signal decay model parameters, generates synthetic images, and uses these for accurate registration [27]. However, these model-based approaches are slower compared to non-iterative methods due to their iterative nature.
@@
Multi-image registration models have been employed by Hanania et al. using a mutual-information-based loss function [12]. Additionally, Gonzales et al. used a multi-scale U-Net to create distance vector fields and applied warping layers to deform feature maps from coarse to fine, which helps to align images with varying intensities by progressively refining the registration and reducing motion artifacts [8, 7]. Li et al. utilized a sparse coding technique to extract the contrast component from the images and subsequently applied a self-supervised deep neural network (DNN) model for image registration [18]. Another approach by Li et al. used spatial-temporal and physical constraints to enforce low-rank and sparsity, with a deep neural network for efficient T1 mapping, denoising, and artifact reduction [17].
@@
Qin et al. used a joint learning of motion estimation and segmentation to improve registration accuracy by leveraging the unsupervised features learned during motion estimation from a large amount of unannotated data [21]. Pan et al. and Hanania et al. employed GANs to generate virtual images after style transfer at different inversion times from a single input image for use in image registration. The GANs reduce the sensitivity of the models to the inherent intensity differences by standardizing contrast across images, which helps to align images with varying contrasts more accurately [20, 11].
@@
Yang et al. introduced a sequential method for addressing contrast differences in images. This method initially separates intensity changes from different inversion times from the constant anatomical structure. However, it heavily relies on accurately disentangling the anatomical structure from the contrast. Additionally, the registration process only takes place between the isolated anatomical images, disregarding the signal decay model along the inversion time axis. Consequently, this approach could result in physically unrealistic deformations due to the necessity for direct integration with the signal decay model [31].
@@
All of these deep learning-based methods, however, do not directly consider the signal decay model, which is essential as the image contrast changes according to it. As a result, these methods may produce physically unrealistic deformations. While classical model-based approaches address these issues, their iterative nature makes them quite slow.
@@
Previously we introduced PCMC-T1, a method specifically designed for free-breathing myocardial T1 mapping that incorporates physical constraints into the motion correction process [13]. This deep learning-based approach ensures accurate T1 mapping under free-breathing conditions by utilizing a supervised framework. However, the practical application of PCMC-T1 is limited by its reliance on large annotated datasets and extensive training, which are not readily available for every scanner and protocol.
@@
In this work, we introduce a self-supervised approach extending our previous PCMC-T1 model. The proposed MBSS-T1 approach enables the generation of corrected T1 maps across various imaging protocols, eliminating the need for protocol-specific data collection, manual annotations, and extensive training.
@@ 3 Methods
$M_{Z}(M_{0},T1,t)=M_{0}\cdot\left(1-2\cdot e^{-t/T1}\right)$ (1)
@@
This classical least-squares approach, as well as deep learning methods such as the one described in [9], assumes that the images are motion-free, typically achieved through prior image registration techniques. However, in practice, motion can be a significant factor, especially in cardiac imaging. To account for motion, the optimization problem must be extended to include deformation fields, which can be introduced as additional variables in the least-squares problem.
@@
$\widehat{T1},\widehat{M_{0}},\widehat{\Phi}=\mathop{\mathrm{arg\,min}}\nolimits_{T1,M_{0},\Phi}\sum_{i=0}^{N-1}\left\|M_{Z}(M_{0},T1,t)_{i}-\phi_{i}\circ I_{i}\right\|^{2}$ (3)
@@
To overcome these limitations, we reframe the optimization task as estimating the weights of a deep neural network ($\Theta$) that simultaneously predicts the T1 model parameters and spatial transformations from the observed magnetization signals and corresponding time points. This is achieved by using two encoder-decoder modules that operate concurrently: one for predicting the T1 parameters (parametric mapping component) and one for predicting the deformation fields (motion component). The objective function to be minimized is expressed as:
$\Theta=\mathop{\mathrm{arg\,min}}\nolimits_{\Theta}\sum_{i=0}^{N-1}\left\|M_{Z}(f_{\Theta}[1]_{M_{0}},f_{\Theta}[1]_{T1},t)_{i}-f_{\Theta}[0]_{i}\circ I_{i}\right\|^{2}$ (4)
@@ 3.1 Model Architecture
Figure 2: Overview of the MBSS-T1 model architecture: (a) The motion component is responsible for deformable image registration, predicting deformation fields that align the acquired images across different time points. (b) The parametric mapping component estimates the T1 and M0 parameters and generates synthetic aligned T1-weighted (T1W) images. (c) The segmentation and uncertainty component uses a pre-trained segmentation network to extract the myocardium segmentations, calculates the deformed segmentations, and retains the ones with high confidence. The network aims to minimize the distance between the motion-corrected T1W images and the synthetic images while maximizing the Dice score between the deformed segmentations.
Fig.2 presents our overall architecture. Our network comprises two U-Net-inspired encoder-decoder modules that operate simultaneously. The network has three main components:
Motion Estimation Component: The upper module, an extension of the VoxelMorph architecture, functions as a multi-image deformable image registration module. It dynamically adjusts the number of registration heads based on the number of recovery inversion time points ($N$) and outputs a specific deformation field $\{\phi_{i}\mid i=0,\ldots,N-1\}$ and registered image $\{R_{i}\mid i=0,\ldots,N-1\}$ for each timestamp $i$. This module leverages the deformation field prediction layers from [2] and the integration layer from [3].
@@
Both encoder-decoder modules follow a U-Net-like structure and include convolutional and downsampling layers, as well as skip connections that establish links between the encoder and decoder of each module.
Segmentation and Uncertainty Component: To ensure the anatomical correctness of the aligned images, we added anatomical constraints to the registration process. First, nnU-Net [16] (other segmentation networks can also be used) is utilized as a pretrained network to predict the myocardium in every unregistered frame. Second, an uncertainty block is employed to retain only the segmentations with high confidence in their correctness. This block receives the masks and softmax maps generated by the nnU-Net output and retains only those masks that meet two criteria: (1) a required percentage $\gamma$ of the myocardial segmentation pixels must have a score greater than a confidence threshold $\alpha$, and (2) the segmentation mask must consist of exactly one connected component. Here, $P_{i}(x)$ denotes the softmax probability at pixel $x$ in the $i$-th segmentation, and $S_{i}$ is the corresponding segmentation mask. Let $G$ be the set of indices of segmentations that satisfy both of these criteria:
$G=\left\{i\mid\frac{1}{|S_{i}|}\sum_{x\in S_{i}}\mathbf{1}\left[P_{i}(x)>\alpha\right]\geq\gamma\text{ and }S_{i}\text{ is connected}\right\}$ (5)
Next, the masks are aligned using the deformation fields $\Phi=\{\phi_{i}\}_{i=0}^{N-1}$ from the Motion Estimation Component to obtain deformed segmentations. During the registration process, the Dice score is maximized between the deformed myocardium segmentations to ensure that the myocardiums are properly registered. We define a reference image index $r$ as the time point corresponding to the segmentation with the highest confidence score:
$r=\underset{i\in G}{\arg\max}\left(\frac{1}{|S_{i}|}\sum_{x\in S_{i}}P_{i}(x)\right)$ (6)

### v2

@@ Abstract
T1 mapping is a valuable quantitative MRI technique for diagnosing diffuse myocardial diseases. Traditional methods, relying on breath-hold sequences and echo triggering, face challenges with patient compliance and arrhythmias, limiting their effectiveness. Image registration can enable motion-robust T1 mapping, but inherent intensity differences between time points pose a challenge. We introduce MBSS-T1, a self-supervised model for motion correction in cardiac T1 mapping, constrained by physical and anatomical principles. The physical constraints ensure expected signal decay behavior, while the anatomical constraints maintain realistic deformations. The unique combination of these constraints ensures accurate T1 mapping along the longitudinal relaxation axis. MBSS-T1 outperformed baseline deep-learning-based image registration approaches in a 5-fold experiment on a public dataset of 210 patients (STONE sequence) and an internal dataset of 19 patients (MOLLI sequence). MBSS-T1 excelled in model fitting quality ($R^{2}$: 0.975 vs. 0.941, 0.946), anatomical alignment (Dice score: 0.89 vs. 0.84, 0.88), and expert visual quality assessment for the presence of visible motion artifacts (4.33 vs. 3.38, 3.66). MBSS-T1 has the potential to enable motion-robust T1 mapping for a broader range of patients, overcoming challenges such as arrhythmias and suboptimal compliance, and allowing for free-breathing T1 mapping without requiring large training datasets. Our code will be publicly available upon acceptance.
@@ 1 Introduction
Quantitative T1 mapping is an advanced MRI technique designed to precisely measure intrinsic longitudinal relaxation time in myocardial tissue [27]. This method is increasingly recognized as essential for assessing diffuse myocardial diseases such as inflammation, fibrosis, hypertrophy, and infiltration [27, 29]. “Native” T1 mapping, performed without a paramagnetic contrast agent, effectively detects myocardial conditions, including edema, iron overload, infarcts, and scarring [25]. Furthermore, by allowing for the objective quantification of tissue characteristics, T1 mapping enables longitudinal tracking of changes in the myocardium. This feature is precious for clinical trials, where monitoring the efficacy of interventions over time is critical [25].
@@
Figure 1 presents the process of generating accurate T1 maps. It begins with capturing a series of images at different time points. Each image captures the myocardium at a distinct point during the longitudinal relaxation. These images are then analyzed to calculate T1 relaxation times through signal decay model fitting, resulting in detailed T1 maps that reflect the myocardial tissue properties. However, heart motion, respiration, and spontaneous patient movements can cause significant distortions in T1 maps, affecting their reliability and clinical usefulness and potentially leading to incorrect diagnoses [28]. Echo-triggering is a well-established method for accounting for cardiac motion. Breath-hold sequences, such as the Modified Look-Locker Inversion Recovery (MOLLI) sequence and its variants [24], are often used to minimize respiratory motion artifacts. However, requiring patients to hold their breath presents practical challenges, as not all patients can fully adhere to these instructions. This breath-hold acquisition further limits the number of images that can be captured and makes T1 mapping unsuitable for patients who cannot tolerate breath-holding [24]. Additionally, inadequate echo-triggering due to cardiac arrhythmias can lead to unreliable T1 maps, compromising diagnostic accuracy.
@@
Despite advancements in classical and deep learning-based methods for motion correction and T1 mapping, significant limitations persist. Classical model-based approaches, while effective, are often slow due to their iterative nature [28, 31]. On the other hand, deep-learning-based methods rely heavily on extensively annotated datasets, which are not readily available for all scanners and protocols [8, 32]. Furthermore, these methods typically use either pairwise loss functions based on metrics such as mutual information to guide the registration process [1] or groupwise approaches such as the principal component analysis (PCA), which ignore the signal decay relaxation model [34]. This omission may result in physically unlikely deformations due to changes in image contrast. These limitations hinder the practical applicability and robustness of current methods.
This work presents MBSS-T1, a self-supervised, physically, and anatomically constrained deep learning model for simultaneous motion correction and T1 mapping from either breath-hold or free-breathing acquisitions. Our method integrates physical and anatomical insights to correct motion in T1 mapping. The main innovation lies in incorporating the signal decay model within the network’s architecture in a self-supervised framework while utilizing a pretrained segmentation network to produce segmentations that guide the registration and ensure the anatomical correctness of the deformed images. This approach enables the generation of corrected T1 maps across various imaging protocols, eliminating the need for protocol-specific data collection, manual annotations, and extensive training. This work extends our previous supervised PCMC-T1 model [13], which required a large annotated dataset and extensive training, thereby limiting its generalization capacity across different acquisition settings, such as sequence type and the number of frames acquired.
We have demonstrated the added value of MBSS-T1 in free-breathing cardiac T1 mapping using a 5-fold experimental setup. This was done on a publicly available free-breathing quantitative dataset of 210 patients [4], acquired using the STONE sequence [30], and an in-house dataset of 19 patients acquired using the MOLLI sequence [24], for both free-breathing and breath-hold scans. Our approach was compared to baseline methods for deep-learning-based image registration [3, 15], as well as the T1 maps produced on the scanner. We used the fit quality (r-squared), the Dice coefficient, and the Hausdorff distance as quantitative metrics to assess registration accuracy and clinical usability. We further demonstrated the clinical impact through qualitative expert radiologist assessments.
@@ 2 Related Work
Guyader et al. introduced a groupwise image registration technique for quantitative MRI that utilizes a dissimilarity metric derived from an approximated form of total correlation. This method addresses the challenge of aligning multiple images simultaneously without relying on a reference image, thereby reducing registration bias and improving the consistency of quantitative measurements [10]. Tao et al. introduced a robust motion correction method for myocardial T1 and extracellular volume mapping using a PCA-based groupwise image registration technique, which registers all MOLLI frames simultaneously with a total-correlation-based metric [26]. Similarly, Huizinga et al. developed a PCA-based groupwise image registration technique for quantitative MRI (qMRI), including T1 mapping, to correct misalignment without a reference image, thus avoiding registration bias. Their method registers all images to a mean space by minimizing a PCA-derived cost function, assuming that a low-dimensional signal model can represent intensity changes in qMRI without needing details of the specific acquisition model [16]. Segmentation-based registration techniques address intensity differences between images using anatomical structures to guide the registration. El-Rewaidy et al. utilized a non-rigid active shape model-based framework, registering T1-weighted images and computing residual motion by matching manually annotated myocardium contours. This method improves T1 map quality by addressing motion-induced errors in myocardial segments [4]. Hering et al. further advanced cardiac motion tracking by combining global semantic information from segmentation labels with local distance metrics, improving the alignment of surrounding structures [14].
A key limitation of the methods above is their failure to account for the signal decay model during registration, resulting in potentially non-physical corrected T1 maps. Model-based approaches have been developed to address this, significantly enhancing T1 mapping accuracy. Xue et al. proposed a fast variational non-rigid registration framework using synthetic image estimation to correct motion in myocardial T1 mapping. Motion-free synthetic images, constructed from initial T1 estimates, guide the registration to avoid motion artifacts [31]. Van Heeswijk et al. introduced a model-based alignment of Look-Locker MRI sequences for myocardial scar quantification, using exponential curve fitting errors as the registration metric [6]. Tilborghs et al. applied a robust motion correction method for cardiac T1 and ECV mapping, integrating data-driven initialization with model-based registration within a non-rigid framework. This method iteratively estimates signal decay model parameters, generates synthetic images, and uses these for accurate registration [28]. However, their iterative nature makes these model-based approaches slower than non-iterative methods.
@@
Multi-image registration models have been employed by Hanania et al. using a mutual-information-based loss function [12]. Additionally, Gonzales et al. used a multi-scale U-Net to create distance vector fields and applied warping layers to deform feature maps from coarse to fine, which helps to align images with varying intensities by progressively refining the registration and reducing motion artifacts [8, 7]. Li et al. utilized a sparse coding technique to extract the contrast component from the images. Subsequently, they applied a self-supervised deep neural network (DNN) model for image registration [19]. Another approach by Li et al. used spatial-temporal and physical constraints to enforce low-rank and sparsity, with a deep neural network for efficient T1 mapping, denoising, and artifact reduction [18].
@@
Qin et al. used a joint learning of motion estimation and segmentation to improve registration accuracy by leveraging the unsupervised features learned during motion estimation from a large amount of unannotated data [22]. Pan et al. and Hanania et al. employed GANs to generate virtual images after style transfer at different inversion times from a single input image for image registration. The GANs reduce the sensitivity of the models to the inherent intensity differences by standardizing contrast across images, which helps to align images with varying contrasts more accurately [21, 11].
@@
Yang et al. introduced a sequential method for addressing contrast differences in images. This method initially separates intensity changes from different inversion times from the constant anatomical structure. However, it heavily relies on accurately disentangling the anatomical structure from the contrast. Additionally, the registration process only occurs between the isolated anatomical images, disregarding the signal decay model along the inversion time axis. Consequently, this approach could result in physically unrealistic deformations due to the necessity for direct integration with the signal decay model [32].
@@
These deep learning-based methods, however, do not directly consider the signal decay model, which is essential as the image contrast changes according to it. As a result, these methods may produce physically unrealistic deformations. While classical model-based approaches address these issues, their iterative nature makes them relatively slow.
@@
Previously, we introduced PCMC-T1, a method designed explicitly for free-breathing myocardial T1 mapping that incorporates physical constraints into the motion correction process [13]. This deep learning-based approach uses a supervised framework to ensure accurate T1 mapping under free-breathing conditions. However, the practical application of PCMC-T1 is limited by its reliance on large annotated datasets and extensive training, which are not readily available for every scanner and protocol.
@@
This work introduces a self-supervised approach extending our previous PCMC-T1 model. The proposed MBSS-T1 approach enables the generation of corrected T1 maps across various imaging protocols, eliminating the need for protocol-specific data collection, manual annotations, and extensive training.
@@ 3 Methods
$M_{Z}(M_{0},T1,t)=M_{0}\cdot\left(1-2\cdot e^{-t/T1}\right)\,.$ (1)
@@
This classical least-squares approach, as well as deep learning methods such as the one described in [9], assumes that the images are motion-free, typically achieved through prior image registration techniques. However, motion can be a significant factor in practice, especially in cardiac imaging. To account for motion, the optimization problem must be extended to include deformation fields, which can be introduced as additional variables in the least-squares problem.
@@
$\widehat{T1},\widehat{M_{0}},\widehat{\Phi}=\mathop{\mathrm{arg\,min}}\nolimits_{T1,M_{0},\Phi}\sum_{i=0}^{N-1}\left\|M_{Z}(M_{0},T1,t)_{i}-\phi_{i}\circ I_{i}\right\|^{2}\,.$ (3)
@@
To overcome these limitations, we reframe the optimization task as estimating the weights of a deep neural network ($\Theta$) that simultaneously predicts the T1 model parameters and spatial transformations from the observed magnetization signals and corresponding time points. This is achieved using two encoder-decoder modules that operate concurrently: one for predicting the T1 parameters (parametric mapping component) and one for predicting the deformation fields (motion component). The objective function to be minimized is expressed as:
$\Theta=\mathop{\mathrm{arg\,min}}\nolimits_{\Theta}\sum_{i=0}^{N-1}\left\|M_{Z}(f_{\Theta}[1]_{M_{0}},f_{\Theta}[1]_{T1},t)_{i}-f_{\Theta}[0]_{i}\circ I_{i}\right\|^{2}\,.$ (4)
@@ 3.1 Model Architecture
Figure 2: Overview of the MBSS-T1 model architecture: (a) The motion component is responsible for deformable image registration, predicting deformation fields that align the acquired images across different time points. (b) The parametric mapping component estimates the T1 and M0 parameters and generates synthetic aligned T1-weighted (T1W) images. (c) The segmentation and uncertainty component uses a pre-trained segmentation network to extract the myocardium segmentations, calculates the deformed segmentations, and retains the ones with high confidence. The network aims to minimize the distance between the motion-corrected T1W and synthetic images while maximizing the Dice score between the deformed segmentations.
Figure 2 presents our overall architecture. Our network comprises two U-Net-inspired encoder-decoder modules that operate simultaneously. The network has three main components:
Motion Estimation Component: The upper module, an extension of the VoxelMorph architecture, functions as a multi-image deformable image registration module. It dynamically adjusts the number of registration heads based on the number of recovery inversion time points ($N$). It outputs a specific deformation field $\{\phi_{i}\mid i=0,\ldots,N-1\}$ and registered image $\{R_{i}\mid i=0,\ldots,N-1\}$ for each timestamp $i$. This module leverages the deformation field prediction layers from [2], and the integration layer from [3].
@@
Both encoder-decoder modules follow a U-Net-like structure, including convolutional and downsampling layers and skip connections that establish links between each module’s encoder and decoder.
Segmentation and Uncertainty Component: To ensure the anatomical correctness of the aligned images, we added anatomical constraints to the registration process. First, nnU-Net [17] (other segmentation networks can also be used) is utilized as a pretrained network to predict the myocardium in every unregistered frame. Second, an uncertainty block is employed to retain only the segmentations with high confidence in their correctness. This block receives the masks and softmax maps generated by the nnU-Net output. It retains only those masks that meet two criteria: (1) a required percentage $\gamma$ of the myocardial segmentation pixels must have a score greater than a confidence threshold $\alpha$, and (2) the segmentation mask must consist of exactly one connected component. Here, $P_{i}(x)$ denotes the softmax probability at pixel $x$ in the $i$-th segmentation, and $S_{i}$ is the corresponding segmentation mask. Let $G$ be the set of indices of segmentations that satisfy both of these criteria:
$G=\left\{i\mid\frac{1}{|S_{i}|}\sum_{x\in S_{i}}\mathbf{1}\left[P_{i}(x)>\alpha\right]\geq\gamma\text{ and }S_{i}\text{ is connected}\right\}\,.$ (5)
Next, the masks are aligned using the deformation fields $\Phi=\{\phi_{i}\}_{i=0}^{N-1}$ from the Motion Estimation Component to obtain deformed segmentations. During registration, the Dice score is maximized between the deformed myocardium segmentations to ensure that the myocardiums are properly registered. We define a reference image index $r$ as the time point corresponding to the segmentation with the highest confidence score:
$r=\underset{i\in G}{\arg\max}\left(\frac{1}{|S_{i}|}\sum_{x\in S_{i}}P_{i}(x)\right)\,.$ (6)